OpenStudy (anonymous):

If f(x) = 6x – 4, evaluate f(x)/-4 for x = 8. A.–2 B.–11 C.44 D.11

OpenStudy (yrelhan4):

Find the value of f(8).. just subtitute 8 in place of x in f(x).. whatever you get divide that by -4. the value you get is your answer. thats what your question is asking you.

OpenStudy (anonymous):

So Would It Be -2/1?

OpenStudy (anonymous):

As In A. ?

OpenStudy (yrelhan4):

Nope. put 8 in f(x).. what do you get?

OpenStudy (anonymous):

If I Put 8 In f(x) Then It Would Be 8/-4 Which Would Be -2?

OpenStudy (yrelhan4):

when i say put x in f(x) i mean replace the value of x by 8 in f(x).. as your f(x) is 6x - 4.. so you'll get f(8) = 6*8 - 4.. yeah?

OpenStudy (anonymous):

Ok I Understand So It'll Be C

OpenStudy (anonymous):

So What Is The Fraction About?

OpenStudy (yrelhan4):

hold up man. your question also asks you to divide it by -4.. once you get f(8).. divide that value by -4.. that'll be your answer.

OpenStudy (anonymous):

Okaaay . So -11 ! Thanks!

OpenStudy (yrelhan4):

Yeah. You're welcome.

OpenStudy (anonymous):

Well What Would I Do For If f(x) = 6x – 9, evaluate f(x)/3 for x = –5.?

OpenStudy (yrelhan4):

No. Same procedure. Put x=-5 in f(x) to find f(-5) and divide that by 3..

OpenStudy (anonymous):

Yes Thats What I Meant So , 39/3 . And I'd Get 13

OpenStudy (yrelhan4):

Yup. Well done.

OpenStudy (anonymous):

Thanks !

OpenStudy (yrelhan4):

Wait.. -39 for f(-5) and divide that by 3.. You'll get -13. not 13.